OpenStudy (anonymous):

Evaluating Functions, Simplify 9.) f(x)= 1/(sqrt x-1) f(x)-f(2)/x-2

OpenStudy (anonymous):

so what is f(2)?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

So plug 2 into the first equation to find f(2): f(2) = 1/(sqrt 2 - 1) Then plug this into the second expression: [f(x) - f(2)]/[x-2] = [(1/(sqrt x -1)) - (1/(sqrt 2 - 1))]/(x-2) Then simplify.

OpenStudy (anonymous):

\[\frac{f(x)-f(2) }{ x-2 } = \frac{ \frac{ 1 }{ \sqrt{x} -1 } - \frac{ 1 }{ \sqrt{2} -1} }{ x-2 }\]

OpenStudy (anonymous):

okay i got to 1/sqrt -1 - sqrt1 all over x-2

OpenStudy (triciaal):